OpenStudy (anonymous):

How many solutions are in this equation 64x^2-16x+1=0

OpenStudy (dtan5457):

1 solution, which is 1/8 or .125.

OpenStudy (dtan5457):

Simply factor this using the borrow payback method.

OpenStudy (dtan5457):

64x^2-16x+1=0

OpenStudy (dtan5457):

Take out the 64 to multiply with c.

OpenStudy (dtan5457):

You get x^2-16x+64. You then factor this to (x-8)(x-8) Plug back the 64. (64x-8)(64x-8) Same solution, divide by 8. (8x-1)=1/8

OpenStudy (dtan5457):

Clear?